40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  PSL  Not logged in ELOG logo
Entry  Thu Mar 15 14:25:05 2018, Craig, DailyProgress, Other, Cavity Power fluctuations vs Temp Fluctuations Screen_Shot_2018-03-15_at_2.53.16_PM.png
    Reply  Thu Mar 15 15:58:02 2018, Craig, DailyProgress, BEAT, Cavity Power fluctuations vs Temp Fluctuations Beatnote_ASD_gpstime_1205189250.pdf
    Reply  Thu Mar 15 18:37:06 2018, Craig, DailyProgress, Other, Cavity Power fluctuations vs Temp Fluctuations NitrogenCylinderRegulatorBefore.jpgWallAirGaugeBefore.jpgNitrogenCylinderRegulatorAfter.jpgScreen_Shot_2018-03-15_at_6.43.59_PM.png
       Reply  Thu Mar 15 22:48:21 2018, Craig, DailyProgress, Other, Cavity Power fluctuations vs Temp Fluctuations 
       Reply  Mon Mar 19 12:10:27 2018, Craig, DailyProgress, Other, Cavity Power fluctuations vs Temp Fluctuations AirSpringPressureRegulator_20180319.jpg
          Reply  Mon Mar 19 13:29:24 2018, awade, DailyProgress, Other, Cavity Power fluctuations vs Temp Fluctuations 
             Reply  Mon Mar 19 14:22:57 2018, Craig, DailyProgress, Other, Cavity Power fluctuations vs Temp Fluctuations NorthAirSpringTubing2.jpgNorthAirSpringTubing.jpg
                Reply  Fri Mar 30 20:12:22 2018, awade, DailyProgress, Other, Disabling the air springs on the vac can: Shimming and clamping 2018-03-30_17.24.30.jpg2018-03-30_17.24.24.jpg
Message ID: 2137     Entry time: Thu Mar 15 15:58:02 2018     In reply to: 2136
Author: Craig 
Type: DailyProgress 
Category: BEAT 
Subject: Cavity Power fluctuations vs Temp Fluctuations 

Btw when I propped up the vaccan using shims in preparation for venting the air springs, our old friend in the beatnote ASD reappeared: the broad hump from 100-10000 Hz. 
This was a problem for us in Dec-Jan, but it went away and we really didn't understand why at the time.  Turns out it's probably upconversion of seismic activity coupling into our cavities.

Quote:

Our initial impression of cavity power fluctuations was that temperature fluctuations in the EAOMs were the cause.  To check this, I made some REFL DC monitors yesterday.
Plotted is one hour of data from today.  Some notes:
1) TRANS DC and REFL DC for both cavites are breathing together every fifteen minutes, and are anticorrelated (one goes up, the other goes down). 
2) The temperature monitors are not fluctuating with the same regularity as the power monitors.
3) The REFL DC for the north PMC is fluctuating with temperature

This leads me to believe our power fluctuations are caused by changes of alignment into the cavities from the air springs holding up our vacuum can. 
Right now, the air springs are hooked up to the wall.  There is probably some pressure regulator which switches on and off every fifteen minutes.  To fix this, I'm going to switch the vacuum air springs over to our nitrogen cylinder we have in the lab and see if the fluctuations go away.

 

Attachment 1: Beatnote_ASD_gpstime_1205189250.pdf  180 kB  Uploaded Thu Mar 15 17:02:18 2018  | Hide | Hide all
Beatnote_ASD_gpstime_1205189250.pdf
ELOG V3.1.3-